# Self-Supervised Learning in Electron Microscopy: Towards a Foundation Model for Advanced Image Analysis

### v1

@@ 3.3 Downstream Tasks
Semantic segmentation involves labeling each pixel of an image with a corresponding class of what is being represented. In materials and biological sciences, this imaging task plays a crucial role in analyzing and understanding complex microstructural and elemental features within the images obtained. Sytwu et al. conducted experiments on semantic segmentation of Gold nanoparticles of different sizes and resolutions using the U-Net model (Ronneberger et al., 2015). They used a U-Net with different numbers of residual blocks and different receptive field sizes to study the influence of model complexity and receptive field size on the performance of the model. Their findings show that increasing the receptive field increases model performance in high-resolution TEM images. They also conclude that as the model complexity increases, there is a corresponding improvement in performance and prediction confidence. Here, the focus is on an investigation of the influence of pretraining on the performance of DL models in semantic segmentation. Therefore, we selected the same network as that used by Sytwu et al., i.e., the U-Net, and used it as the generator for pretraining on unlabeled data in the previous step. We then fine-tuned it for semantic segmentation on the same dataset, i.e., high-resolution TEM images of Gold nanoparticles. The results were compared to training the same model with randomly initialized weights. These experiments were conducted with different complexities and receptive field sizes. We additionally used a more complex High-Resolution Network (HRNet) (Sun et al., 2019) as the generator and fine-tuned it on this dataset for comparison.
Furthermore, atom segmentation, localization, noise reduction, and deblurring are crucial tasks in atomic-resolution scanning transmission electron microscopy (STEM). The images captured at the atomic scale often suffer from noise, which can obscure subtle details and compromise the accuracy of atom segmentation and localization (Lin et al., 2021). Denoising is important in refining these images, ensuring a clearer representation of the atomic structure. It enhances the level of detail in the images beyond the inherent resolution of the microscope, achieving higher spatial resolution, which allows researchers to discern finer structural features and better characterize atomic arrangements. By reducing noise and enhancing resolution, the method ensures a more accurate and robust analysis of atomic structures, even in challenging imaging conditions with variations in sample thickness. To show the generalizability of pretrained models in our work, we use the TEMImageNet data and do experiments on denoising, noise & background removal, and super-resolution. We use the HRNet model and run the same experiments to compare the results of fine-tuning to those of training with random weight initialization. Details of experiments and results are outlined in Section 4.
@@ 4.1 Experiments on CEM500K and GANs
The CEM500K dataset was employed in the pretraining step where cGANs, based on the Pix2Pix model (Isola et al., 2017), were used. As generator, we used different U-Net architectures with varying numbers of residual blocks and receptive field sizes. Specifically, configurations with 2, 3, and 4 residual blocks were considered, each associated with different receptive field sizes. For two residual blocks, the receptive field sizes were 44, 84, and 116, for three blocks, we used 96, 176, and 240 as well as a receptive field of 200, 360, and 424 for four blocks, in line with the work presented in (Sytwu et al., 2022). For brevity, we refer to these U-Net variations as U-Net_B_RF from now on, where B refers to the number of residual blocks and RF refers to receptive field size. Additionally, a more complex model architecture, HRNet (Wang et al., 2021), was used to explore its effectiveness in comparison to U-Net variants and, in particular, to understand how far an increase in model complexity results in an increase in expressivity.
For training, random Gaussian noise, blurring, flipping, and rotations were applied to the images used as input to the generator. The generator was trained to predict images indistinguishable from the original “clean” images. The training for each model variation was conducted for epochs with a batch size of . Adam optimizer with a learning rate of was used for optimization. In terms of training the whole GAN architecture, it was framed as the Least Square GAN (LSGAN), which adopts the least squares loss function for the discriminator and is more stable than regular GANs. LSGANs are able to generate higher quality images and perform more stably during the learning process (Mao et al., 2017). The generator was also trained using the loss, and the in Equation 3 was set to .
@@
The validation plots for the GAN pretraining with HRNet and the most complex U-Net variation are shown in Figure 4. For each of the models, as the dataset size increases, the validation loss decreases. Interestingly, the U-Net model shows better validation results compared to HRNet. This is also illustrated in the generated images in Figure 5. We believe this is due to the skip connections from the initial residual blocks in the U-Net model to the corresponding upsampling blocks. The U-Net model has direct access to the information in the larger spatial resolutions and is more prone to memorizing, while the HRNet model actually learns the feature representations as it encodes the images into lower spatial resolutions with higher feature maps and then decodes it, without having direct access to the input features in the higher spatial resolution. Additionally, we find that the pretrained HRNet model fine-tuned on a variety of supervised tasks outperforms the U-Net model. For brevity, the results for other U-Net variations and for the other dataset sizes can be found in the Appendix.
@@ 4.2 Experiments on Semantic segmentation with Au datasets
Conducting the same experiments with the HRNet model, we find that a fine-tuned U-Net model still outperforms the randomly initialized HRNet model, but all three fine-tuned HRNet models not only outperform the randomly initialized HRNet model but also score higher than the randomly initialized U-Net, as well as all fine-tuned U-Net models. The validation plots are shown in Figure 9. Example predictions are illustrated in Figure 10 for the fine-tuned HRNet and fine-tuned U-Net_4_0, pretrained on CEM500K images.
The general trend of the results and observations for all other datasets is consistent with the above-reported ones.
@@ 4.3 Experiments on TEMImageNet dataset
The plots for validation loss in all three tasks are shown in Figures 11. We observe that in all three cases, the validation losses for the fine-tuned models are lower than those for the randomly initialized models. During the initial training phase, the validation loss for the model fine-tuned with smaller datasets is larger than that for the models trained with more data. However, already after approximately 10 epochs, this difference vanishes. Additionally, the validation loss of all randomly initialized models exhibits severe fluctuations, while the fine-tuned models behave much more robustly. Some example predictions for denoising, noise & background removal, and super-resolution are shown in Figures 12, 13, and 14, respectively. All predictions are visually very good and it is difficult to observe any differences.
@@ 5 Conclusion
Our exploration of pretraining on unlabeled data, followed by fine-tuning on labeled data, enhanced model performance and accelerated convergence across several downstream tasks, including segmentation, denoising, and super-resolution. An important outcome of our work for such tasks is that for obtaining a higher predictive accuracy, the model complexity might not be the only or most important factor. This has a particular relevance in the context of significantly larger foundation models, for which the present work can be considered as an important stepping stone.
Appendix A Appendix
The appendix here includes example illustrations, tables and formulas that were either directly referred to in the main text or moved here due to page limits.
A.1 Objective Function for cGANs
@@
A.2 Example Predictions for Semantic Segmentation
A.3 Validation Plots for the Pretraining
A.4 Quantitative Evaluations for Semantic Segmentation
@@
A.5 Quantitative Evaluation on TEMImageNet datasets

### v2

@@ 3.3 Downstream Tasks
Semantic segmentation involves labeling each pixel of an image with a corresponding class of what is being represented. In materials and biological sciences, this imaging task plays a crucial role in analyzing and understanding complex microstructural and elemental features within the images obtained. Sytwu et al. conducted experiments on semantic segmentation of Gold nanoparticles of different sizes and resolutions using the U-Net model [37]. They used a U-Net with different numbers of residual blocks and different receptive field sizes to study the influence of model complexity and receptive field size on the performance of the model. Their findings show that increasing the receptive field increases model performance in high-resolution TEM images. They also conclude that as the model complexity increases, there is a corresponding improvement in performance and prediction confidence. Here, the focus is on an investigation of the influence of pretraining on the performance of DL models in semantic segmentation. Therefore, we selected the same network as that used by Sytwu et al., i.e., the U-Net, and used it as the generator for pretraining on unlabeled data in the previous step. We then fine-tuned it for semantic segmentation on the same dataset, i.e., high-resolution TEM images of Gold nanoparticles. The results were compared to training the same model with randomly initialized weights. These experiments were conducted with different complexities and receptive field sizes. We additionally used a more complex High-Resolution Network (HRNet) [45] as the generator and fine-tuned it on this dataset for comparison.
Furthermore, atom segmentation, localization, noise reduction, and deblurring are crucial tasks in atomic-resolution scanning transmission electron microscopy (STEM). The images captured at the atomic scale often suffer from noise, which can obscure subtle details and compromise the accuracy of atom segmentation and localization [26]. Denoising is important in refining these images, ensuring a clearer representation of the atomic structure. It enhances the level of detail in the images beyond the inherent resolution of the microscope, achieving higher spatial resolution, which allows researchers to discern finer structural features and better characterize atomic arrangements. By reducing noise and enhancing resolution, the method ensures a more accurate and robust analysis of atomic structures, even in challenging imaging conditions with variations in sample thickness. To show the generalizability of pretrained models in our work, we use the TEMImageNet data and do experiments on denoising, noise & background removal, and super-resolution. We use the HRNet model and run the same experiments to compare the results of fine-tuning to those of training with random weight initialization. Details of experiments and results are outlined in Section 4.
@@ 4.1 Experiments on CEM500K and GANs
The CEM500K dataset was employed in the pretraining step where cGANs, based on the Pix2Pix model [18], were used. As generator, we used different U-Net architectures with varying numbers of residual blocks and receptive field sizes. Specifically, configurations with 2, 3, and 4 residual blocks were considered, each associated with different receptive field sizes. For two residual blocks, the receptive field sizes were 44, 84, and 116, for three blocks, we used 96, 176, and 240 as well as a receptive field of 200, 360, and 424 for four blocks, in line with the work presented in [46]. For brevity, we refer to these U-Net variations as U-Net_B_RF from now on, where B refers to the number of residual blocks and RF refers to receptive field size. Additionally, a more complex model architecture, HRNet [50], was used to explore its effectiveness in comparison to U-Net variants and, in particular, to understand how far an increase in model complexity results in an increase in expressivity.
For training, random Gaussian noise, blurring, flipping, and rotations were applied to the images used as input to the generator. The generator was trained to predict images indistinguishable from the original “clean” images. The training for each model variation was conducted for epochs with a batch size of . Adam optimizer with a learning rate of was used for optimization. In terms of training the whole GAN architecture, it was framed as the Least Square GAN (LSGAN), which adopts the least squares loss function for the discriminator and is more stable than regular GANs. LSGANs are able to generate higher quality images and perform more stably during the learning process [28]. The generator was also trained using the loss, and the in Equation 3 was set to .
@@
The validation plots for the GAN pretraining with HRNet and the most complex U-Net variation are shown in Figure 4. For each of the models, as the dataset size increases, the validation loss decreases. Interestingly, the U-Net model shows better validation results compared to HRNet. This is also illustrated in the generated images in Figure 5. We believe this is due to the skip connections from the initial residual blocks in the U-Net model to the corresponding upsampling blocks. The U-Net model has direct access to the information in the larger spatial resolutions and is more prone to memorizing, while the HRNet model actually learns the feature representations as it encodes the images into lower spatial resolutions with higher feature maps and then decodes it, without having direct access to the input features in the higher spatial resolution. Additionally, we find that the pretrained HRNet model fine-tuned on a variety of supervised tasks outperforms the U-Net model. For brevity, results for other experiments are included in the supplementary materials.
@@ 4.2 Experiments on Semantic segmentation with Au datasets
Conducting the same experiments with the HRNet model, we find that a fine-tuned U-Net model still outperforms the randomly initialized HRNet model, but all three fine-tuned HRNet models not only outperform the randomly initialized HRNet model but also score higher than the randomly initialized U-Net, as well as all fine-tuned U-Net models. The validation plots are shown in Figure 9. Example predictions are illustrated in Figure 10 for the fine-tuned HRNet and fine-tuned U-Net_4_0, pretrained on CEM500K images. Moreover, the dice scores on test set are calculated for all models after 5, 30 and 60 epochs. As shown in Table 2, the pretrained models converge faster than randomly initialized models and generally score higher, specially in the beginning epochs. The general trend of the results and observations for all other datasets is consistent with the above-reported ones.
| Epochs | 5 | 30 | 60 | ||
| HRNet | R | 41.05 | 86.55 | 90.24 | |
| P(50k) | 86.37 | 91 | 91.97 | ||
| P(100k) | 88.49 | 92.14 | 92.38 | ||
| P(200k) | 83.86 | 91.41 | 92.07 | ||
| U-Net 4 blocks | 424 | R | 7.8 | 86.03 | 89.95 |
| P(50k) | 72.78 | 89.81 | 91.16 | ||
| P(100k) | 71.8 | 89 | 89.72 | ||
| P(200k) | 75.46 | 87.34 | 90.21 | ||
| 200 | R | 0.25 | 80.69 | 83.61 | |
| P(50k) | 82.03 | 87.05 | 85.68 | ||
| P(100k) | 83 | 86.7 | 88.53 | ||
| P(200k) | 55.75 | 86.56 | 89 | ||
| U-Net 2 blocks | 116 | R | 0 | 80.81 | 82.49 |
| P(50k) | 61.99 | 82.3 | 84.72 | ||
| P(100k) | 51.97 | 80.16 | 83.47 | ||
| P(200k) | 67.87 | 80.48 | 83.18 | ||
| 44 | R | 0 | 74.87 | 79.75 | |
| P(50k) | 65.89 | 76.9 | 78.59 | ||
| P(100k) | 58.62 | 75.67 | 79.92 | ||
| P(200k) | 53.15 | 75.45 | 76.99 | ||
@@ 4.3 Experiments on TEMImageNet dataset
The plots for validation loss in all three tasks are shown in Figure 11. We observe that in all three cases, the validation losses for the fine-tuned models are lower than those for the randomly initialized models. During the initial training phase, the validation loss for the model fine-tuned with smaller datasets is larger than that for the models trained with more data. However, already after approximately 10 epochs, this difference vanishes. Additionally, the validation loss of all randomly initialized models exhibits severe fluctuations, while the fine-tuned models behave much more robustly. Some example predictions for denoising, noise & background removal, and super-resolution are shown in Figures 12, 13, and 14, respectively. Even though all predictions are visually very good and it is difficult to observe any differences, the quantitative results as confirmed by the plots in Figure 11 show that pretraining leads to a better and more stable performance. Similarly, the metric on test set are calculated for all models after 5, 30 and 60 epochs. As shown in Table 3, the pretrained models converge faster than randomly intialized models and generally score better, specially in the beginning epochs.
| Epochs | 5 | 30 | 60 | |
| SR | R | 0.01972 | 0.01524 | 0.01122 |
| P(50k) | 0.01688 | 0.01359 | 0.0112 | |
| P(100k) | 0.02105 | 0.01465 | 0.01142 | |
| P(200k) | 0.01639 | 0.01371 | 0.01156 | |
| N&BGR | R | 0.06391 | 0.03438 | 0.0176 |
| P(50k) | 0.03332 | 0.02257 | 0.01776 | |
| P(100k) | 0.03268 | 0.02913 | 0.01766 | |
| P(200k) | 0.03323 | 0.0224 | 0.01792 | |
| DN | R | 0.04732 | 0.02996 | 0.01541 |
| P(50k) | 0.03287 | 0.02169 | 0.01542 | |
| P(100k) | 0.03693 | 0.03708 | 0.01555 | |
| P(200k) | 0.02657 | 0.01907 | 0.01586 | |
@@ 5 Conclusion
Pretraining on unlabeled data, followed by fine-tuning on labeled data, enhanced performance and accelerated convergence in several downstream tasks, including segmentation, denoising, and super-resolution. An important outcome of our work for such tasks is that for obtaining a higher predictive accuracy, the model complexity might not be the only or most important factor. The CEM500K dataset, containing SEM images, was used for pretraining and improved performance in TEM image-based downstream tasks, despite the differences between SEM and TEM images. Future research could explore pretraining on TEM images for closer domain relevance. Additionally, while self-supervised learning with GANs enhanced the performance in this study, their training complexities and risk of mode collapse suggest exploring alternative self-supervised methods like contrastive pretraining for potentially better outcomes in the future.
Supplementary Material for Self-Supervised Learning with Generative Adversarial Networks for Electron Microscopy
Bashir Kazimi1, Karina Ruzaeva1, Stefan Sandfeld1,2
1Forschungszentrum Jülich GmbH, Jülich, IAS-9, Germany
2RWTH Aachen University, Aachen, Germany
{b.kazimi,k.ruzaeva,s.sandfeld}@fz-juelich.de
This Supplementary Material includes example illustrations, tables and formulas that were either directly referred to in the main text or moved here due to page limits.
5.1 Objective Function for cGANs
@@
5.2 Example Predictions for Semantic Segmentation
5.3 Validation Plots for the Pretraining
5.4 Quantitative Evaluations for Semantic Segmentation
@@
5.5 Quantitative Evaluation on TEMImageNet datasets